OpenStudy (mikey):

fan and medal Solve 9x2 + 6 = 12x. 1 plus or minus i square root 10 the quantity of 2 plus or minus i square root 10 all over 3 the quantity of 2 plus or minus i square root 2 all over 3 1 plus or minus i square root 2

OpenStudy (studygurl14):

Put into this format: \(\large ax^2+bx+c=0\) Then use the quadratic formula \(\Huge\frac{-b\pm\sqrt{b^2-4ac}}{2a}\)

OpenStudy (mikey):

i dont understand that formula at all, SORRY

OpenStudy (studygurl14):

You have \(\large 9x^2 +6=12x\) right?

OpenStudy (mikey):

right

OpenStudy (studygurl14):

Okay, so rearrange to fit the standard form \(\large ax^2+bx+c=0\) --> \(\large 9x^2 -12x + 6=0\) With me so far?

OpenStudy (mikey):

yes im along

OpenStudy (studygurl14):

Okay, so notice the form of \(\large ax^2+bx+c=0\) and \(\large 9x^2-12x+6=0\) match. Therefore, you can see that a = 9, b = -12, and c= 6 Understand?

OpenStudy (mikey):

yes i just dont understand the second equation you gave me

OpenStudy (studygurl14):

Okay, see those variables and values I gave you. Just plug those into the formula, and simplify. That will get you the value of x

OpenStudy (studygurl14):

\(\Huge\frac{-(-12)\pm\sqrt{(-12)^2-4(9)(6)}}{2(9)}\)

OpenStudy (mikey):

so the b^2 will be -12

OpenStudy (mikey):

ah okay

OpenStudy (studygurl14):

Let me know if you need any other help

OpenStudy (mikey):

thank you soso very much

OpenStudy (studygurl14):

You're welcome :)